If the first and third of consecutive even integers are added, the result is 22 less than three times the second integer. Find the integers
Found 2 solutions by mananth, HasanSahin:
Answer by mananth(16946)

If the first and third of consecutive even integers are added, the result is 22 less than three times the second integer. Find the integers
let x,x+1,x+2 be the integers
...
x+x+2=3(x+1)-22
2x+2=3x+3-22
3x-2x=22-3+2
x=21
...
the integers are 21,22,23
..
m.ananth@hotmail.ca

Answer by HasanSahin(52)

x,x+2,x+4 these are the consecutive even integers..
x+x+4 = 2x+4 = 3*(x+2)-22 >> 2x-4 = 3x+6-22 >> x=12 >>12,14,16 are the integers
RF.